OpenStudy (henrietepurina):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (kc_kennylau):

y=-x-4: slope=-1, y-intercept=-4

OpenStudy (kc_kennylau):

y-1=-1(x-(-3))

OpenStudy (kc_kennylau):

y-1=-x-3

OpenStudy (kc_kennylau):

y=-x-2

OpenStudy (kc_kennylau):

B